# Floating-Base Deep Lagrangian Networks

### v1

@@ IV-E First mass moment
where $\mathbf{K}(\mathbf{q})=[\mathbf{L}_{1\mathrm{L}}(\mathbf{q}_{1}),\dots,\mathbf{L}_{{\mathrm{n}_{\mathrm{k}}}\mathrm{L}}(\mathbf{q}_{\mathrm{n}_{\mathrm{k}}})]$. Therefore, the constraint (20) can be fulfilled by expressing $\mathbf{L}_{\mathrm{R}\mathrm{L}}(\mathbf{q})$ in terms of $\mathbf{L}_{\mathrm{R}}$, $\mathbf{K}$, $\mathbf{W}$, and the skew-symmetric matrix $\mathbf{S}_{\mathbf{h}}$,
@@ IV-F Mass
where $\mathbf{U}=[\mathbf{L}_{\mathrm{R}\mathrm{L}},\mathbf{K}]$. Thus, we express $\mathbf{L}_{\mathrm{L}}$ in terms of the total mass $\mathrm{m}>0$ and $\mathbf{U}$ as
@@ IV-H Parametrization Overview
(${\mathrm{n}_{\mathrm{q}}}=12$) (${\mathrm{n}_{\mathrm{q}}}=17$) (${\mathrm{n}_{\mathrm{q}}}=22$)
$\mathbf{H}$ 324 529 784
Standard $\mathbf{C}$ 171 276 406
Reordered $\mathbf{L}$ 117 162 210
Proposed 106 151 199
Body Parameters 208 288 368
V Floating-Base DeLaN
@@ V-A Inertia Matrix
To enforce input-independencies for each kinematic branch, we define $\mathrm{n}_{\mathrm{k}}$} networks $\mathbf{NN}_{k}$ to estimate $\mathbf{L}_{k\mathrm{L}}(\mathbf{q}_{k})$, $\mathbf{L}_{k\mathrm{R}}(\mathbf{q}_{k})$ and $\mathbf{L}_{k}(\mathbf{q}_{k})$ using only the joints $\mathbf{q}_{k}$ of the $k$th branch as input. Meanwhile, $\mathbf{L}_{\mathrm{R}}(\mathbf{q})$ and $\mathbf{L}_{\Sigma}(\mathbf{q})$ are estimated by a single network $\mathbf{NN}_{\mathrm{R}}$ that takes all joint positions as input, while $\mathrm{m}$ is computed by $\theta_{\mathrm{m}}^{2}$ which is an unconstrained trainable parameter.
@@
where $\mu_{\mathbf{D}}$ is the smallest eigenvalue of $\mathbf{D}$, and $\epsilon_{\mathrm{D}}$ is a small offset. Analogously, we ensure $\mathbf{T}\succ 0$ in (24) using
$\hat{\mathrm{m}}$ $=\mathrm{softplus}(\mathrm{m}-\lambda_{\mathbf{U}})+\epsilon_{\mathrm{m}}+\lambda_{\mathbf{U}}\text{,}$ (27)
@@
where $\lambda_{\mathbf{U}}$ is the largest eigenvalue of the product $\mathbf{U}^{\mathrm{\scriptscriptstyle T}}\mathbf{U}$ and $\epsilon_{\mathrm{m}}$ is a small offset. Both $\mathbf{U}^{\mathrm{\scriptscriptstyle T}}\mathbf{U}$ and $\mathbf{W}^{\mathrm{\scriptscriptstyle T}}\mathbf{W}$ can be interpreted as the contribution of all the kinematic branches to the composite spatial inertia. Therefore, $\epsilon_{\mathrm{m}}$ and $\epsilon_{\mathrm{D}}$ can be interpreted as minimal bounds regarding the contribution of the robot’s base to the spatial inertia. Furthermore, $\hat{\mathrm{m}}$ and $\hat{\mathbf{D}}$ do not affect any of the physical consistency constraints. Finally, Algorithm 1 summarizes the inertia matrix computation from the networks’ outputs and (21), (23), (27), (28).
@@
0: joint position $\mathbf{q}$, parameters $\theta_{\mathrm{m}}$, $\boldsymbol{\phi}_{\mathrm{R}},\dots,\boldsymbol{\phi}_{\mathrm{n}_{\mathrm{k}}}]$
@@ V-D Loss Function
The Lagrangian (1) cannot be learned directly with standard supervised learning, as the system’s energy is not observed. Following the approach of [36], the energies are learned indirectly by minimizing the error from the estimated torque $f^{-1}(\boldsymbol{\nu},\dot{\boldsymbol{\nu}},\ddot{\boldsymbol{\nu}},\boldsymbol{\phi})$ given by Algorithm 2. More precisely, by optimizing the following problem
@@
$+w_{\mathbf{U}}(\hat{\mathrm{m}}-\mathrm{m})^{2}+w_{\mathbf{D}}\mathrm{softplus}(-\mu_{\mathbf{D}})^{2}\text{,}$
where $\mathbf{W}_{\boldsymbol{\tau}}$ is the diagonal covariance of $\boldsymbol{\tau}_{\nu}$ in the training set; $w_{\mathbf{U}}$ and $w_{\mathbf{D}}$ are weights for auxiliaries terms from (25) and (26). Note that physical consistency is achieved as a hard constraint due to the structured parametrization, the auxiliary terms only encourage the intermediate values $\mathrm{m}$ and $\mathbf{D}$ to be positive and positive-definite, respectively, which empirically accelerated training during our experiments. We also apply the input layer $\mathbf{T}_{\mathbf{q}}(\mathbf{q})=[\cos(\mathbf{q}),\sin(\mathbf{q})]$ to each network, as proposed in [36].
Algorithm 2 Torque Computation
@@
3:$P=$ (31) with $\mathrm{m}$, $\mathbf{h}(\mathbf{q})$
@@ VI Experiments
VI-A Baselines
To evaluate the trade-off of physics assumptions and estimation accuracy, we compare our method against different baselines for estimating the inverse dynamics model. We employ a feed-forward neural network (FFNN) with inputs $\boldsymbol{\Theta},\mathbf{q},\dot{\boldsymbol{\nu}},\ddot{\boldsymbol{\nu}}$. Additionally, we use DeLaN [36] with one network using $\mathbf{q}$ as input to estimate $\mathbf{C}(\mathbf{q})$ of $\mathbf{H}(\mathbf{q})$ in (32), and a second network that uses the entire $\boldsymbol{\nu}$ to estimate $P(\boldsymbol{\nu})$. To evaluate the proposed potential energy parametrization in (31), we implement a modified version of DeLaN, namely DeLaN-PP, which uses a single network for $\mathbf{H}(\mathbf{q})$ and computes $P$ from its estimated terms using (31). We also consider a variant of our proposed network FeLaN, named FeLaN-BS, that incorporates only branch-induced sparsity and positive definiteness. In this variant, $\mathbf{L}_{\mathrm{L}},\mathbf{L}_{\mathrm{R}\mathrm{L}},$ and $\mathbf{L}_{\mathrm{R}}$ are estimated by a neural network. We add the input layer $\mathbf{T}_{\mathbf{q}}(\mathbf{q})$ to every entry, with the exception of the FFNN and the linear coordinates in DeLaN, which do not receive this input. We use 2 hidden layers of 32 neurons for FFNN, and 2 layers of 16 neurons for the other networks.
Regarding white-box methods, we use Mujoco XLA [38] as a differentiable recursive Newton-Euler (DNEA) function parametrized by 16 kinodynamic body parameters. In this case, physical consistency is connected to the inertial parameters. Therefore, we evaluate parameterization with different degrees of physical consistency:
  • DNEA NS [20]: unconstrained inertial parameters
  • DNEA Tri [20]: (6) with a geometric parametrization
All methods were implemented and trained in JAX [40] using stochastic gradient descent. Training details are provided in Appendix A.
VI-B Simulated Robots
For an initial evaluation, with robots mainly governed by rigid-body dynamics without actuator dynamics, we collect data at 100 Hz for Go2 and Talos in MuJoCo XLA [38]. For Talos, we consider only the first four arm joints, assuming the remaining joints as fixed. To ensure sufficient system excitation and data diversity, we command uniformly sampled desired velocities and base heights to be tracked by the model predictive controller proposed in [41] for both robots. For Talos’ arms, we provide sine-wave reference trajectories with uniformly sampled amplitudes and frequencies. Each dataset yields a total of 50k samples. As the ground-truth model is available, we compute $\boldsymbol{\tau}_{\nu}$ using (3). Table II presents the normalized mean squared error (NMSE) of the estimated torques computed with $\mathbf{W}_{\boldsymbol{\tau}}$ for all methods. To compare performance across robots, we define the relative NMSE as $r\text{NMSE}_{i}=(\text{NMSE}_{i}-\text{NMSE}_{\min})/\text{NMSE}_{\max}$ for every method in the test set (also shown in Tab. II).
For Go2, the physically consistent DNEAs achieved the best performance, since the estimated parameters capture the full dynamics. Among the grey-box methods, FeLaN-BS showed similar performance, while the remaining grey-box methods were satisfactory. For Talos, FeLaN-BS and FeLaN achieved errors close to DNEAs, even outperforming DNEA Cov and DNEA Log. Although the amplitude errors remain small, the slightly worse performance of DNEA on Talos may be due to contacts adding larger spikes in the data when compared to Go2, which make the learning more challenging. For both robots, the FFNN performed worst. Nevertheless, FeLaN-BS achieved the best $r\text{NMSE}$.
TABLE II: Inverse dynamics normalized mean squared error for simulated Go2 and Talos.
Method Go2 Talos $r\text{NMSE}$
Train Test Train Test
FFNN $9.5\mathrm{e}{\text{-}1}\pm 4.3\mathrm{e}{\text{-}2}$ $1.4\mathrm{e}{0}\pm 9.3\mathrm{e}{\text{-}2}$ $5.3\mathrm{e}{0}\pm 2.8\mathrm{e}{\text{-}1}$ $5.0\mathrm{e}{0}\pm 2.1\mathrm{e}{\text{-}1}$ $9.9\mathrm{e}{\text{-}1}$
DeLaN $2.6\mathrm{e}{\text{-}2}\pm 1.0\mathrm{e}{\text{-}2}$ $3.1\mathrm{e}{\text{-}2}\pm 1.3\mathrm{e}{\text{-}2}$ $3.1\mathrm{e}{\text{-}1}\pm 1.1\mathrm{e}{\text{-}1}$ $2.8\mathrm{e}{\text{-}1}\pm 8.0\mathrm{e}{\text{-}2}$ $3.4\mathrm{e}{\text{-}2}$
DeLaN-$P$ $2.5\mathrm{e}{\text{-}2}\pm 6.9\mathrm{e}{\text{-}3}$ $3.3\mathrm{e}{\text{-}2}\pm 2.0\mathrm{e}{\text{-}2}$ $9.0\mathrm{e}{\text{-}1}\pm 1.3\mathrm{e}{0}$ $2\mathrm{e}{\text{-}1}\pm 1.1\mathrm{e}{0} 8.8$$8\mathrm{e}{\text{-}2}$
FeLaN-BS $1.7\mathrm{e}{\text{-}3}\pm 5.4\mathrm{e}{\text{-}4}$ $3.0\mathrm{e}{\text{-}3}\pm 5.1\mathrm{e}{\text{-}3}$ $\mathbf{5.9\mathrm{e}{\text{-}2}\pm 1.3\mathrm{e}{\text{-}2}}$ $\mathbf{5.4\mathrm{e}{\text{-}2}\pm 1.2\mathrm{e}{\text{-}2}}$ $\mathbf{8.7\mathrm{e}{\text{-}4}}$
FeLaN $4.3\mathrm{e}{\text{-}2}\pm 1.9\mathrm{e}{\text{-}3}$ $5.9\mathrm{e}{\text{-}2}\pm 3.5\mathrm{e}{\text{-}3}$ $3.1\mathrm{e}{\text{-}1}\pm 1.2\mathrm{e}{0}$ $2.7\mathrm{e}{\text{-}1}\pm 1.1\mathrm{e}{0}$ $4.4\mathrm{e}{\text{-}2}$
DNEA NS $1.3\mathrm{e}{\text{-}2}\pm 1.9\mathrm{e}{\text{-}2}$ $1.2\mathrm{e}{\text{-}2}\pm 1.8\mathrm{e}{\text{-}2}$ $2.3\mathrm{e}{\text{-}1}\pm 6.3\mathrm{e}{\text{-}1}$ $2.0\mathrm{e}{\text{-}1}\pm 5.5\mathrm{e}{\text{-}1}$ $1.8\mathrm{e}{\text{-}2}$
DNEA PD $6.8\mathrm{e}{\text{-}4}\pm 6.2\mathrm{e}{\text{-}5}$ $6.3\mathrm{e}{\text{-}4}\pm 1.3\mathrm{e}{\text{-}4}$ $8\mathrm{e}{\text{-}1}\pm 7.0\mathrm{e}{\text{-}1} 2.2$$2\mathrm{e}{\text{-}1}\pm 5.6\mathrm{e}{\text{-}1}$ $1.6\mathrm{e}{\text{-}2}$
DNEA Tri $6.8\mathrm{e}{\text{-}4}\pm 6.1\mathrm{e}{\text{-}5}$ $6.3\mathrm{e}{\text{-}4}\pm 1.1\mathrm{e}{\text{-}4}$ $3.9\mathrm{e}{\text{-}1}\pm 6.5\mathrm{e}{\text{-}1}$ $2.9\mathrm{e}{\text{-}1}\pm 4.6\mathrm{e}{\text{-}1}$ $2.4\mathrm{e}{\text{-}2}$
DNEA Cov $6.8\mathrm{e}{\text{-}4}\pm 4.7\mathrm{e}{\text{-}5}$ $6.2\mathrm{e}{\text{-}4}\pm 1.2\mathrm{e}{\text{-}4}$ $3.7\mathrm{e}{\text{-}1}\pm 1.1\mathrm{e}{0}$ $3.1\mathrm{e}{\text{-}1}\pm 9.0\mathrm{e}{\text{-}1}$ $2.6\mathrm{e}{\text{-}2}$
DNEA Log $\mathbf{6.7\mathrm{e}{\text{-}4}\pm 5.0\mathrm{e}{\text{-}5}}$ $\mathbf{6.1\mathrm{e}{\text{-}4}\pm 8.4\mathrm{e}{\text{-}5}}$ $5.6\mathrm{e}{\text{-}1}\pm 1.0\mathrm{e}{0}$ $4.7\mathrm{e}{\text{-}1}\pm 8.4\mathrm{e}{\text{-}1}$ $4.1\mathrm{e}{\text{-}2}$
Figure 1 shows each component in (3) of the linear force $z$ and two Talos joint torques. All grey-box methods approximate the total torque well (see Fig. 1(a)), but only FeLaN preserves a stationary gravity component for $z$ due to proper mass parametrization.
@@ VI-B Simulated Robots
Figure 1: Estimated force $z$ and torques for left arm (LA) second joint, and right leg (RL) fifth joint for simulated Talos.

### v2

@@ IV-E First mass moment
where $\mathbf{K}(\mathbf{q})=[\mathbf{L}_{1\mathrm{L}}(\mathbf{q}_{1});\dots;\mathbf{L}_{{\mathrm{n}_{\mathrm{k}}}\mathrm{L}}(\mathbf{q}_{\mathrm{n}_{\mathrm{k}}})]$. Therefore, the constraint (20) can be fulfilled by expressing $\mathbf{L}_{\mathrm{R}\mathrm{L}}(\mathbf{q})$ in terms of $\mathbf{L}_{\mathrm{R}}$, $\mathbf{K}$, $\mathbf{W}$, and the skew-symmetric matrix $\mathbf{S}_{\mathbf{h}}$,
@@ IV-F Mass
where $\mathbf{U}=[\mathbf{L}_{\mathrm{R}\mathrm{L}};\mathbf{K}]$. Thus, we express $\mathbf{L}_{\mathrm{L}}$ in terms of the total mass $\mathrm{m}>0$ and $\mathbf{U}$ as
@@ IV-H Parametrization Overview
(${\mathrm{n}_{\mathrm{q}}}=12$) (${\mathrm{n}_{\mathrm{q}}}=19$) (${\mathrm{n}_{\mathrm{q}}}=22$)
$\mathbf{H}$ 324 625 784
Standard $\mathbf{C}$ 171 325 406
Reordered $\mathbf{L}$ 117 187 210
Proposed 106 176 199
Body Parameters 208 320 368
V Floating-Base Deep Lagrangian Networks
@@ V-A Inertia Matrix
To enforce input-independencies for each kinematic branch, we define $\mathrm{n}_{\mathrm{k}}$} networks $\mathbf{NN}_{k}$ to estimate $\mathbf{L}_{k\mathrm{L}}$, $\mathbf{L}_{k\mathrm{R}}$, and $\mathbf{L}_{k}$ only from the joints $\mathbf{q}_{k}$ of the $k$th branch. Meanwhile, $\mathbf{L}_{\mathrm{R}}$ and $\mathbf{L}_{\Sigma}$ are estimated by a single network $\mathbf{NN}_{\mathrm{R}}$ that takes all joints $\mathbf{q}$ as input, while $\mathrm{m}$ is computed by $\theta_{\mathrm{m}}^{2}$, where $\theta_{\mathrm{m}}\in\mathbb{R}$ is an unconstrained trainable scalar.
@@
where $\mu_{\mathbf{D}}$ is the smallest eigenvalue of $\mathbf{D}$, and $\epsilon_{\mathrm{D}}$ is a small offset. Analogously, we ensure $\mathbf{T}\succ 0$ in (24) by
$\hat{\mathrm{m}}$ $=\mathrm{softplus}(\mathrm{m}-\max(\boldsymbol{\lambda}_{\mathbf{U}}))+\epsilon_{\mathrm{m}}+\max(\boldsymbol{\lambda}_{\mathbf{U}})\text{,}$ (27)
@@
where $\boldsymbol{\lambda}_{\mathbf{U}}$ are the eigenvalues of the product $\mathbf{U}^{\mathrm{\scriptscriptstyle T}}\mathbf{U}$ and $\epsilon_{\mathrm{m}}$ is a small offset. Both $\mathbf{U}^{\mathrm{\scriptscriptstyle T}}\mathbf{U}$ and $\mathbf{W}^{\mathrm{\scriptscriptstyle T}}\mathbf{W}$ can be interpreted as the contribution of all the kinematic branches to the composite spatial inertia. Therefore, $\epsilon_{\mathrm{m}}$ and $\epsilon_{\mathrm{D}}$ can be interpreted as minimal bounds regarding the contribution of the robot’s base to the spatial inertia. Furthermore, $\hat{\mathrm{m}}$ and $\hat{\mathbf{D}}$ do not affect any of the physical consistency constraints. Finally, Algorithm 1 summarizes the inertia matrix computation from the networks’ outputs and (21), (23), (27), (28).
@@
0: joint position $\mathbf{q}$, parameters $(\theta_{\mathrm{m}}$, $\boldsymbol{\phi}_{\mathrm{R}},\dots,\boldsymbol{\phi}_{\mathrm{n}_{\mathrm{k}}})$
@@ V-D Loss Function
TABLE II: Inverse dynamics NMSE over 10 seeds for simulated Go2 and Talos.
Method Go2 Talos $r\text{NMSE}$
Train Test Train Test
MLP $7.5\mathrm{e}{\text{-}2}\pm 1.1\mathrm{e}{\text{-}2}$ $1.1\mathrm{e}{\text{-}1}\pm 3.1\mathrm{e}{\text{-}2}$ $1.9\mathrm{e}{0}\pm 1.8\mathrm{e}{\text{-}1}$ $3.9\mathrm{e}{0}\pm 4.9\mathrm{e}{\text{-}1}$ $9.9\mathrm{e}{\text{-}1}$
DeLaN $2.2\mathrm{e}{\text{-}3}\pm 1.4\mathrm{e}{\text{-}3}$ $4.2\mathrm{e}{\text{-}3}\pm 6.3\mathrm{e}{\text{-}3}$ $1.7\mathrm{e}{\text{-}1}\pm 1.3\mathrm{e}{\text{-}2}$ $2.8\mathrm{e}{\text{-}1}\pm 4.7\mathrm{e}{\text{-}2}$ $4.4\mathrm{e}{\text{-}2}$
DeLaN-$P$ $4.0\mathrm{e}{\text{-}3}\pm 1.9\mathrm{e}{\text{-}3}$ $6.9\mathrm{e}{\text{-}3}\pm 6.9\mathrm{e}{\text{-}3}$ $8.6\mathrm{e}{\text{-}1}\pm 1.4\mathrm{e}{0}$ $1.1\mathrm{e}{0}\pm 1.5\mathrm{e}{0}$ $1.7\mathrm{e}{\text{-}1}$
FeLaN-BS $1.2\mathrm{e}{\text{-}3}\pm 5.5\mathrm{e}{\text{-}4}$ $1.9\mathrm{e}{\text{-}3}\pm 2.8\mathrm{e}{\text{-}3}$ $5.1\mathrm{e}{\text{-}2}\pm 1.0\mathrm{e}{\text{-}2}$ $1.2\mathrm{e}{\text{-}1}\pm 1.5\mathrm{e}{\text{-}1}$ $1.3\mathrm{e}{\text{-}2}$
FeLaN $1.4\mathrm{e}{\text{-}3}\pm 5.5\mathrm{e}{\text{-}4}$ $1.7\mathrm{e}{\text{-}3}\pm 1.2\mathrm{e}{\text{-}3}$ $\mathbf{3.3\mathrm{e}{\text{-}2}\pm 1.1\mathrm{e}{\text{-}2}}$ $\mathbf{6.7\mathrm{e}{\text{-}2}\pm 3.6\mathrm{e}{\text{-}2}}$ $\mathbf{4.9\mathrm{e}{\text{-}3}}$
DNEA NS $3.2\mathrm{e}{\text{-}2}\pm 5.7\mathrm{e}{\text{-}2}$ $2.9\mathrm{e}{\text{-}2}\pm 5.0\mathrm{e}{\text{-}2}$ $3.8\mathrm{e}{\text{-}1}\pm 6.2\mathrm{e}{\text{-}1}$ $4.7\mathrm{e}{\text{-}1}\pm 6.7\mathrm{e}{\text{-}1}$ $1.8\mathrm{e}{\text{-}1}$
DNEA PD $1.7\mathrm{e}{\text{-}3}\pm 4.7\mathrm{e}{\text{-}3}$ $1.4\mathrm{e}{\text{-}3}\pm 3.5\mathrm{e}{\text{-}3}$ $3.6\mathrm{e}{\text{-}1}\pm 7.0\mathrm{e}{\text{-}1}$ $4.3\mathrm{e}{\text{-}1}\pm 7.6\mathrm{e}{\text{-}1}$ $5.0\mathrm{e}{\text{-}2}$
DNEA Tri $9.5\mathrm{e}{\text{-}4}\pm 1.6\mathrm{e}{\text{-}3}$ $8.0\mathrm{e}{\text{-}4}\pm 1.1\mathrm{e}{\text{-}3}$ $5.5\mathrm{e}{\text{-}1}\pm 9.8\mathrm{e}{\text{-}1}$ $6.8\mathrm{e}{\text{-}1}\pm 1.1\mathrm{e}{0}$ $8.0\mathrm{e}{\text{-}2}$
DNEA Cov $\mathbf{6.8\mathrm{e}{\text{-}4}\pm 4.9\mathrm{e}{\text{-}5}}$ $6.4\mathrm{e}{\text{-}4}\pm 2.6\mathrm{e}{\text{-}4}$ $2.5\mathrm{e}{\text{-}1}\pm 6.8\mathrm{e}{\text{-}1}$ $2.8\mathrm{e}{\text{-}1}\pm 7.1\mathrm{e}{\text{-}1}$ $2.8\mathrm{e}{\text{-}2}$
DNEA Log $6.9\mathrm{e}{\text{-}4}\pm 1.3\mathrm{e}{\text{-}4}$ $\mathbf{6.2\mathrm{e}{\text{-}4}\pm 1.1\mathrm{e}{\text{-}4}}$ $3.0\mathrm{e}{\text{-}1}\pm 8.2\mathrm{e}{\text{-}1}$ $3.1\mathrm{e}{\text{-}1}\pm 7.6\mathrm{e}{\text{-}1}$ $3.2\mathrm{e}{\text{-}2}$
The Lagrangian (1) cannot be learned directly with standard supervised learning, as the system’s energy is not observed. Following [26], the energies are learned indirectly by minimizing the error from the estimated torque $f^{-1}(\boldsymbol{\nu},\dot{\boldsymbol{\nu}},\ddot{\boldsymbol{\nu}},\boldsymbol{\phi})$ given by Algorithm 2, by optimizing
@@
$+w_{\mathbf{U}}{\textstyle\sum}{}^{\scriptscriptstyle 3}_{\scriptscriptstyle i=1}\mathrm{softplus}(\boldsymbol{\lambda}_{\mathbf{U}}{}_{i}-\mathrm{m})+w_{\mathbf{D}}\mathrm{softplus}(-\mu_{\mathbf{D}})^{2}$
where $\mathbf{W}_{\boldsymbol{\tau}}$ is the diagonal covariance of $\boldsymbol{\tau}_{\nu}$ in the training set; $w_{\mathbf{U}}$ and $w_{\mathbf{D}}$ weight the auxiliaries terms from (25) and (26). Note that physical consistency is achieved as a hard constraint due to the structured parametrization, the extra terms only encourage $\mathrm{m}$ and $\mathbf{D}$ to be positive and positive-definite, respectively, which empirically accelerated training during our experiments. We also apply the input layer $\mathbf{T}_{\mathbf{q}}(\mathbf{q})=[\cos(\mathbf{q});\sin(\mathbf{q})]$ to each network, as in [26].
Algorithm 2 Inverse Dynamics
@@
3:$P=$ (31) with $\hat{\mathrm{m}}$, $\mathbf{h}(\mathbf{q})$
@@ VI Experiments
Figure 2: Components of the estimated force $z$ and left arm second joint (LA2) torque for simulated Talos.
VI-A Baselines
To evaluate the trade-off of physics assumptions and estimation accuracy, we compare our method against different baselines for estimating the inverse dynamics model. We employ a multilayer perceptron (MLP) with inputs $\boldsymbol{\Theta}$, $\mathbf{q}$, $\dot{\boldsymbol{\nu}}$, $\ddot{\boldsymbol{\nu}}$. Additionally, we use DeLaN [26] with one network using $\mathbf{q}$ as input to estimate $\mathbf{C}$ in (32), and a second network that uses the entire $\boldsymbol{\nu}$ to estimate $P(\boldsymbol{\nu})$. To evaluate the proposed potential energy parametrization in (31), we implement a modified version of DeLaN, namely DeLaN-PP, which uses a single network for $\mathbf{H}(\mathbf{q})$ and computes $P$ from its estimated terms using (31). We also consider a variant of our proposed network FeLaN, named FeLaN-BS, that incorporates only branch-induced sparsity and positive definiteness. In this variant, $\mathbf{L}_{\mathrm{L}},\mathbf{L}_{\mathrm{R}\mathrm{L}},$ and $\mathbf{L}_{\mathrm{R}}$ are estimated by a single network. We add the input layer $\mathbf{T}_{\mathbf{q}}$ to the inputs of all models, except for MLP and the linear coordinates in DeLaN. We use 2 hidden layers of 32 neurons for MLP, and 2 layers of 16 neurons for the other networks.
Regarding white-box methods, we use Mujoco XLA (MJX) [35] as a differentiable recursive Newton-Euler (DNEA) function, parametrized by 16 kinodynamic parameters per body. In this case, physical consistency is connected to the inertial parameters. Therefore, we evaluate parameterization with different degrees of physical consistency:
  • DNEA NS [34]: unconstrained inertial parameters
  • DNEA Tri [37]: (6) via second moments of mass
All methods are implemented in JAX [6] and trained with stochastic gradient optimization for 2k epochs on quadrupeds and 3k on Talos. We use two hyperparameter sets: one for Talos and one for all the quadrupeds. Results are averaged over 10 seeds. Code and datasets are available online [^1] https://schulze18.github.io/felan_website.
VI-B Simulated Robots
For an initial evaluation on robots mainly governed by rigid-body dynamics, we collect data at 100 Hz for Go2 and Talos in MJX. For Talos, we consider only the first four arm joints, while the remaining joints are kept fixed. To ensure sufficient excitation and data diversity, we uniformly sample desired velocities and base heights and track them with the model predictive controller [2] on both robots. For Talos’ arms, we provide sine-wave reference trajectories with uniformly sampled amplitudes and frequencies. Each dataset yields a total of 50k samples. As the ground-truth model is available, we compute $\boldsymbol{\tau}_{\nu}$ using (3). Table II reports the mean and standard deviation over seeds of the torque mean squared error, normalized by $\mathbf{W}_{\boldsymbol{\tau}}$, for all methods, denoted as NMSE. To compare performance across robots, we define the relative NMSE as $r\text{NMSE}_{i}=(\text{NMSE}_{i}-\text{NMSE}_{\min})/\text{NMSE}_{\max}$ for each method on the test set, also shown in Tab. II.
TABLE III: Inverse dynamics NMSE over 10 seeds for real robot data.
Method HyQReal2 Spot Spot with Arm Talos $r\text{NMSE}$ $\overline{r\text{NMSE}}$
Train Test Train Test Train Test Train Test Sim + Real
MLP $\mathbf{3.0\pm 0.4}$ $\mathbf{4.5\pm 0.4}$ $\mathbf{5.9\pm 0.2}$ $\mathbf{4.2\pm 0.3}$ $\mathbf{7.8\pm 0.4}$ $\mathbf{9.3\pm 0.5}$ $28.0\pm 0.1$ $27.3\pm 0.1$ $0.202$ $0.464$
DeLaN $92.5\pm 3.6$ $213.1\pm 104.2$ $6.9\pm 0.1$} $5.1\pm 0.1$} $17.2\pm 12.9$ $10.7\pm 2.2$ $\mathbf{3.8\pm 0.2}$ $\mathbf{5.2\pm 2.0}$ $0.307$ $0.219$
DeLaN-$P$ $6.9\pm 1.8$} $8.2\pm 2.4$} $8.0\pm 0.2$ $5.9\pm 0.1$ $11.1\pm 0.3$ $9.6\pm 0.2$ $7.5\pm 0.3$ $7.3\pm 0.2$ $\mathbf{0.086}$ $0.114$
FeLaN-BS $7.6\pm 0.4$ $9.0\pm 0.6$ $8.4\pm 0.1$ $6.3\pm 0.1$ $10.5\pm 0.2$} $9.5\pm 0.3$} $8.1\pm 0.2$ $7.8\pm 0.2$ $0.103$ $0.073$
FeLaN $6.6\pm 0.1$ $7.8\pm 0.2$ $8.1\pm 0.1$ $6.0\pm 0.1$ $10.5\pm 0.2$} $9.6\pm 0.2$ $7.5\pm 1.0$ $7.2\pm 0.5$ $0.090$ $\mathbf{0.062}$
DNEA NS $12.8\pm 0.2$ $14.5\pm 0.2$ $10.1\pm 0.1$ $7.6\pm 0.1$ $14.4\pm 0.2$ $10.9\pm 0.2$ $15.9\pm 0.5$ $14.8\pm 0.5$ $0.248$ $0.225$
DNEA PD $12.9\pm 0.2$ $14.7\pm 0.3$ $10.1\pm 0.1$ $7.6\pm 0.1$ $14.4\pm 0.2$ $10.9\pm 0.2$ $17.1\pm 1.0$ $16.0\pm 1.0$ $0.260$ $0.190$
DNEA Tri $13.0\pm 0.2$ $14.8\pm 0.2$ $10.2\pm 0.1$ $7.6\pm 0.1$ $4.5\pm 0.2 11$$1.0\pm 0.2$ $17.5\pm 1.0$ $16.4\pm 0.8$ $0.266$ $0.204$
DNEA Cov $13.0\pm 0.2$ $14.8\pm 0.2$ $10.1\pm 0.1$ $7.6\pm 0.1$ $4.5\pm 0.3 11$$1.0\pm 0.3$ $17.5\pm 1.0$ $16.4\pm 0.8$ $0.266$ $0.186$
DNEA Log $13.2\pm 0.1$ $15.0\pm 0.2$ $10.2\pm 0.1$ $7.6\pm 0.1$ $4.5\pm 0.2 11$$1.0\pm 0.2$ $17.7\pm 1.0$ $16.5\pm 0.9$ $0.268$ $0.189$
For Go2, the fully physically consistent DNEAs achieve the best performance, as their consistent parametrization fully captures the system’s dynamics. All grey-box methods perform well, with FeLaN and FeLaN-BS outperforming DNEA NS, demonstrating the importance of physical consistency. For Talos, FeLaN performed best, followed by FeLaN-BS. Although the amplitude errors remain small, the slightly worse performance of DNEA on Talos may be due to contact changes, adding larger spikes in the data when compared to Go2, making the estimation more challenging. DeLaN still performed similarly compared to the other methods, while DeLaN-PP demonstrated sensitivity to different initialization. For both robots, the MLP performed the worst. Accordingly, FeLaN demonstrated the most consistent results, with the lowest $r\text{NMSE}$, followed by FeLaN-BS.
Figure 2 shows the components in (3) for the force $z$ and one torque joint of Talos. All grey-box methods approximate the total value well, see Fig. 2(a). FeLaN better predicts the individual terms, in particular the stationary gravity component of $z$ in Fig. 2(d), due to its proper mass parametrization.